### v2

@@ 3.1 Unified Audio-Visual Encoding
Audio and visual representations. In order to unify representations of different data to minimize data discrepancies, we use a general model [49] pre-trained by aligning vision, audio and language modalities to extract video representations. Specifically, the visual and audio encoders are utilized to tokenize visual and audio snippets of a given input video, respectively. Since the visual encoder just learned from image data while temporal information in videos is essential for TAL and AVEL tasks, the visual encoder fine-tuned on Kinetics-400 [7] is also considered. After tokenization, the visual $E_{V}=\{e_{t}^{v}\}_{t=1}^{T}\in\mathbb{R}^{T\times D}$ and audio $E_{A}=\{e_{t}^{a}\}_{t=1}^{T}\in\mathbb{R}^{T\times D}$ feature sequences can be obtained, where $D$ is the projected feature dimension. Note that the parameters of the visual and audio encoders are frozen during this process.
@@
Audio-visual pyramid transformer. Audio and visual signals are equally crucial for AVEL [47, 11], and both serve as important complementary information for TAL and SED to boost the performance [1, 5, 15]. Therefore, we uniformly feed the video data with both modalities of all tasks into a single backbone, which is implemented by an audio-visual pyramid transformer inspired by [59, 11]. Specifically, the tokenized visual $E_{V}$ and audio $E_{A}$ sequences added with its position embeddings first pass $L_{1}$ transformer blocks separately for uni-modal long-form temporal modeling. In untrimmed videos, the occurring instances usually have various lengths. Thus, to detect them at multiple temporal resolutions, the obtained features from two modalities are fed into an audio-visual pyramid fusion module, which consists of $L_{2}$ transformer blocks. In each block, 2x downsampling using strided convolutions is first applied, and then audio-visual cross-attention is conducted by assigning the current modality as the key and value vectors while another as the query vector. Afterward, the enhanced features from different modalities in each pyramid level $l$ are concatenated to get an audio-visual feature pyramid $Z=\{Z^{l}\}_{l=1}^{L_{2}}$, where $Z^{l}\in\mathbb{R}^{T_{l}\times 2D}$.
3.2 Task-Specific Experts
@@
Unifying these three tasks under a single framework is inherently challenging as they focus on instances with different characteristics. Inspired by mixture-of-experts (MoE) networks [41, 3, 37], we introduce task-specific experts in our transformer blocks. Unlike previous MoEs that aim to capture modality-specific information (e.g., vision and language), our experts allow the model to learn distinct knowledge for different tasks. Specifically, as shown in the right of Fig. 2, the output feature $F_{l-1},l\in[1,L_{1}+L_{2}]$ from a previous block first passes a shared multi-head attention (MHA) to align information from different tasks. Then, the task-specific information can be captured by switching to different task experts. The task experts compose a Multiway feed-forward network (FFN), where each expert is a standard FFN consisting of two linear layers and an activation.
$F_{l}^{{}^{\prime}}=\mathrm{MHA}(\mathrm{LN}(F_{l-1})+F_{l-1}),\;$ (1)
@@
3.3 Head Design
After the audio-visual encoding in Sec. 3.1, the output features $\{Z^{l}\}_{l=1}^{L_{2}}$ from the audio-visual feature pyramid will connect to the classification and regression heads to get localization predictions in a single pass.
Language-aware classification head. Since the datasets for different tasks have different label vocabularies, it is a straightforward way to use task-specific heads for classification. However, it results in parameter redundancy and lacks flexibility due to fixed categories. Here, we propose to unify three task-specific heads into one by taking advantage of large pre-trained models. In detail, we treat the instance categories of each dataset as text information and encode them using a pre-trained text encoder. To add contextual information, prompts are also customized to help specify labels from different tasks. The used prompt templates are “A visual event of {label}.”, “An audio visual event of {label}.”, and “A sound event of {label}.” for TAL, AVEL and SED, respectively. Note that the text encoder is from the same pre-trained model as the visual and audio encoders in Sec. 3.1, which provides a strong prior on measuring the relevancy between modalities. The texts are encoded as $\mathcal{T}=\{\mathcal{T}_{i}\}_{i=1}^{N_{k}}$, where $N_{k}$ is the number of classes of the dataset for the $k$-th task. Then, the encoded texts are linearly projected to a shared embedding space $D^{\prime}$ with the features from the audio-visual feature pyramid. We obtain the normalized text vector $\hat{\mathcal{T}}=\{\hat{\mathcal{T}_{i}}\}_{i=1}^{N_{k}}\in\mathbb{R}^{N_{k}\times D^{\prime}}$ and the normalized cross-modal feature vector $\hat{Z}^{l}\in\mathbb{R}^{T_{l}\times D^{\prime}}$ from each pyramid level to compute the similarities between them.
@@ 3.3 Head Design
where $s^{l}\in\mathbb{R}^{T_{l}\times N_{k}}$ indicates the similarities between $N_{k}$ categories and $T_{l}$ temporal segments in the pyramid level $l$, and $\sigma$ is a learnable scaling factor used to adaptively adjust the magnitude of the similarities as in [39]. Then, a sigmoid function is attached to predict the probabilities of $N_{k}$ classes at each moment.
Regression head. We simply apply a lightweight regression head for each task. Each regression head as in [11] consists of 3 layers of 1D convolution attached with layer normalization and ReLU activation. The parameters of the first two convolutional layers are shared among the three heads. The last layer followed by ReLU outputs the distances to the start/end time of an instance at each moment in the pyramid level $l$ if the instance exists.
Loss function. Following [59, 11], we apply two losses to train our model in an end-to-end manner. The sigmoid focal loss [28] and the generalized IoU loss [40] are used for classification and regression, respectively. The contributions of these two losses are equal as default.
3.4 Multi-Task Training
We jointly train TAL, AVEL and SED tasks in a single, unified architecture by multi-task learning. The advantage is that the tasks can potentially learn mutually beneficial knowledge from each other. However, the datasets for the tasks vary greatly in size and difficulty, making it very challenging for joint training. For instance, a single epoch of ActivityNet 1.3 [6] for TAL task corresponds to around 15 epochs of DESED [20] for SED task. Besides, all clips in DESED are 10s while the longest video in ActivityNet 1.3 exceeds 12 minutes. Hence, following the multi-task training method [32], we use Round-Robin Batch-Level Sampling strategy to sample batches from tasks in a cyclical manner, where one iteration forwards a batch for each task and updates parameters in sequence. The Dynamic Stop-and-Go training scheduler is also applied to monitor the validation losses of each task to avoid overfitting.
4 Experiments
4.1 Datasets and Evaluation Metrics
@@
4.2 Implementation Details
The sampling rates of sounds and video frames are 16 kHz and 16 fps, respectively. We feed 16 consecutive frames using a sliding window with stride 8 for ActivityNet 1.3, and downsample the features into a fixed length of 256 following [59]. For UnAV-100 and DESED, we use the stride of 4, and pad or crop the feature sequences to 256 and 64, respectively. For each corresponding 1s audio segment, the same stride duration (0.5/0.25s) is used to temporally align with the visual ones. The visual, audio and text encoders of ONE-PEACE [49] are utilized to extract semantically aligned three modality embeddings, where the visual encoder is further fine-tuned on Kinetics-400 [7]. The extracted feature dimension is 1,536 for all three modalities. The dimension of the embedding space in the framework is $D=D^{{}^{\prime}}=512$, and $L_{1}=2$, $L_{2}=6$. During training, we use Adam for optimization and simply set the same hyperparameters for all datasets of three tasks. Specifically, the mini-batch size is 16, the initial learning rate is 1e-3 and a cosine learning rate decay is used. Our model is trained only for 5 epochs with a linear warmup of 2 epochs and the weight decay is 1e-4. During inference, our model outputs the on/offsets and classes with confidence scores for all three types of instances occurring in a given video. The output candidates are then processed by Soft-NMS [4] to eliminate highly overlapping ones.
Table 1: Comparison of our multi-task models to single-task performance. We use both audio and visual modalities for all models. “ST”: single-task, “AT”: all tasks. “All Tasks Average” is computed by averaging the average mAP results of all three tasks. “$\cdot\mathrm{M}(\cdot)$”: total $\cdot\mathrm{M}$ parameters needed for $\cdot$ models to conduct all three tasks.
ActivityNet 1.3 (TAL) UnAV-100 (AVEL) DESED (SED) All Tasks Average
@@ 4.2 Implementation Details
2 Multi-Task (Base) 55.4 34.9 6.3 34.8 53.0 47.4 41.0 33.5 21.5 49.8 62.7 50.5 26.2 58.6 47.7 62M (1)
3 Multi-Task (TAL & AVEL) 56.4 35.7 5.0 35.7 54.0 48.9 41.9 34.0 21.7 50.8 - - - - - 97M (1)
4 Multi-Task (TAL & SED) 56.5 36.1 4.2 35.5 - - - - - - 61.8 49.3 26.1 58.2 - 97M (1)
5 Multi-Task (AVEL & SED) - - - - 51.8 46.8 40.1 30.0 18.2 48.3 62.6 49.4 27.2 59.4 - 97M (1)
6 Multi-Task (AT) 57.1 36.4 5.5 36.1 54.1 48.6 42.1 34.3 20.5 50.7 64.0 49.5 27.0 59.8 48.9 130M (1)
[…]
4.3 Multi-Task Performance
@@
Single-task v.s. multi-task. To demonstrate the effectiveness of our unified framework, in Tab. 1, we compare our multi-task (AT) model (row 6) with two baseline models, single-task (ST) and multi-task (base) in row 1-2. Single-task (ST) models were trained individually on the three tasks, using standard transformer blocks with one FFN and a 1D-conv classifier similar to the regressor. The multi-task (base) model has the same architecture as ST models except for using multi-task learning strategies in Sec. 3.4. We use the same hyperparameters described in Sec. 4.2 as the default setting unless specified otherwise. We observe that simply joint training cannot get decent results on the three tasks (row 2). The performance of the multi-task (base) model on SED and AVEL saw a slight improvement, due to beneficial knowledge learned from each other. However, the results on TAL declined since there exists a significant gap between ActivityNet 1.3 and other datasets. In contrast, our multi-task (AT) model (row 6) that applies the task-specific experts and the unified language-aware classifier can achieve performance boosts on all three tasks compared with single-task models, e.g., $+0.8\%$, $+1.1\%$ and $+2.1\%$ at the average mAP on TAL, AVEL and SED, respectively, and $+1.4\%$ at the average performance of all tasks. Besides, the total number of parameters reduces by a factor of 1.4×1.4$\times$1.4 × (i.e., 186M to 130M), going from 3 full models to only 1 required for all tasks. It implies both the effectiveness and efficiency of our multi-task framework.
Pair-wise task relationships. We also explore pair-wise task relationships by jointly training two of three tasks, shown in rows 3-5 in Tab. 1. We observe that when applying our proposed experts and unified classifier, both AVEL and SED (row 3-4) can benefit from the rich common instances in ActivityNet 1.3 when trained with TAL, leading to an obvious improvement compared with the single-task models, i.e., $+1.2\%$ and $+0.5\%$ at the average mAP, respectively. Besides, SED gains a significant boost when trained with AVEL (row 5), i.e., $+1.7\%$ at the average mAP. It could be attributed to category overlap between UnAV-100 and DESED, and training with the large dataset can help prevent overfitting in the small one. Conversely, due to the large gap in instance duration and dataset scales, SED tends to have a negative effect on the AVEL task, resulting in a decrease of $1.3\%$ at the average mAP (row 5). But this effect can be regulated by jointly training all three tasks together using our proposed model (row 6).
Multi-task learning as pre-training. Inspired by [32], we finetune each single-task model on our trained AT model to demonstrate that the AT model can allow downstream tasks to take advantage of multi-task training. The results are shown in row 8 of Tab. 1, where we initialize ST models using the trained AT model and finetune them individually using the same training recipe as in row 1. We can see that the ST models fine-tuned on our AT model outperform the single-task models in row 1 by a large margin, i.e., $+0.9\%$, $+2.1\%$ and $+3.4\%$ on TAL, AVEL and SED, respectively. It indicates that joint training can capture knowledge that is mutually beneficial to all these three tasks, being an effective pre-training step for single-task models.
4.4 Comparison with Existing Work
Table 2 presents the comparison results of our model with state-of-the-art works. For TAL task, we note that many previous TAL methods achieved superior results on ActivityNet 1.3 by combining with the external action classifier [53]. By contrast, our model has good capabilities in both classification and regression and does not need to rely on any external classification models. For a fair comparison, we list the results of methods without using external classifiers in Tab. 2. We simply concatenate audio and visual features as input for those TAL methods [30, 42, 59, 60] when using both modalities. We can see that, with ONE-PEACE [49] features, our single all-task (AT) model reaches an average mAP of $36.1\%$, outperforming state-of-the-art task-specific models. Here we also trained UnAV [11] that is tailored for AVEL task to conduct TAL task, but found the result is much lower than highly specialized TAL models [42, 59]. It indicates that the task-specific models cannot be generalized effectively to other tasks, while our unified model can achieve superior or on-par performances on all three tasks with good generalizability. For AVEL task, our AT model gets competitive results to UnAV [11] when using the same ONE-PEACE features. And our single-task fine-tuned model (OursAT→ST) further improves the average mAP to $51.7\%$, setting a new state-of-the-art result on the AVEL task. For SED task, since the traditional SED methods only support super fine-grained sound spectrograms as input with the evaluation metric not suitable for our unified approach, we implemented Actionformer [59], TriDet [42] and UnAV [11] to conduct SED task. We can see that our AT model outperforms all other methods by a large margin. Overall, we emphasize that our UniAV holds superior or on-par performances on all three tasks using a single unified model, and achieves the best All Tasks Average score compared with other methods, which significantly distinguishes our model from other task-specific models.
@@ 4.4 Comparison with Existing Work
Table 2: Comparison with existing state-of-the-art methods. We report the mAP at tIoU=0.5 and the average mAP on three tasks. Best results are in bold and second best underlined. “OP-V/A” denotes the visual/audio encoder of ONE-PEACE [49]. “*” denotes that the results for the AVEL task are from UnAV [11].
@@
Method Visual Encoder Audio Encoder ActivityNet 1.3 (TAL) UnAV-100 (AVEL) DESED (SED) All Tasks Average
@@
4.5 Ablation Studies
Effect of task-specific experts. To verify our design choices, we explore the effect of applying task-specific experts on different layers of the pyramid transformer. As shown in Tab. 3, applying task-specific experts brings a significant performance boost. Specifically, using experts in the later $L_{2}$ blocks (row 3) has better results than that in the early $L_{1}$ blocks (row 2), which indicates that the later stages of the model can capture distinct knowledge that is more beneficial for each task. Besides, adding experts in all transformer blocks (row 6) achieves the best performances on two of three tasks (i.e., TAL and SED).
Language-aware classification head. We also perform ablations on our proposed language-aware classification head (LCH) as shown in Tab. 3. We find that using task-specific classifiers (row 4) leads to quite unstable results on three tasks. By contrast, notable improvements are observed on TAL and SED tasks when we use LCH with only categories as text tokens (row 5). Especially for SED, there is a significant $4.5\%$ increase at average mAP. It could be attributed to the unified LCH based on the large language encoder [49], which enhances the model’s generalization, effectively avoiding overfitting on the quite small dataset. Moreover, using prompts to add context information mentioned in Sec. 3.3 can further improve the performance on two of the three tasks.
Audio-visual fusion for TAL and SED. We also verify the effectiveness of audio-visual fusion for TAL and SED tasks in Tab. 4(a). We use the single-task model (row 1 in Tab. 1) for each task to conduct the experiments. For TAL, we can see the performance increase ($+1.1\%$ at average mAP) as we insert audio signals to apply cross-modal interactions. For SED, the model obtains a substantial performance boost ($+8.2\%$ at average mAP) when adding visual modalities, indicating the critical role of both modalities for the task.
Effect of different visual encoders. As shown in Tab. 4(b), we compare the performances of our models using different visual encoders. For both single-task (ST) and multi-task (AT) models, the performances improve by a large margin when using the visual encoder [49] fine-tuned on Kinetics-400 [7]. In particular, for TAL, the improvements of $3.4\%$ and $3.5\%$ at average mAP can be observed for the ST and AT models, respectively. This emphasizes the importance of motion information for video localization tasks, especially for the TAL task.
Table 3: Ablation study on the main proposed components. “E-$L_{1}$” denotes applying task-specific experts on the early $L_{1}$ transformer blocks, and “E-$L_{2}$” denotes applying experts on the later $L_{2}$ blocks. “LCH” is short for language-aware classification head. “Prompt” denotes using prompts when tokenizing instance categories for different tasks.

### v3

@@ III-A Unified Audio-Visual Encoding
Audio and visual generic representations. Individually collected datasets for different tasks exhibit significant variations in domains and scales. In order to unify representations of diverse data to minimize data discrepancies, we use a general model [21] pre-trained by aligning vision, audio and language modalities to extract video representations. Specifically, the visual and audio encoders of the model [21] are utilized to tokenize visual and audio snippets of a given input video, respectively. Note that the parameters of the encoders are frozen during this process. Since the visual encoder just learned from image data while temporal information in videos is essential for TAL and AVEL tasks, the visual encoder fine-tuned on Kinetics-400 [60] is also considered. After tokenization, we apply a linear projection layer to project the audio and visual representations into a shared embedding space. Then, the visual $E_{V}=\{e_{t}^{v}\}_{t=1}^{T}\in\mathbb{R}^{T\times D}$ and audio $E_{A}=\{e_{t}^{a}\}_{t=1}^{T}\in\mathbb{R}^{T\times D}$ embedding sequences can be obtained, where $D$ is the projected feature dimension.
Audio-visual pyramid transformer. We consider that there is often irrelevant information and noise in audio and video tracks that interfere with accurate localization for specific types of instances, such as off-screen voice/music for AVEL, a visible dog without barking for SED. Besides, diverse modality-aware instances occurring in untrimmed videos might span various duration. For instance, the event duration ranges from 0.2 to 60 seconds in UnAV-100 [15] for AVEL. Therefore, an effective cross-modal fusion strategy is essential for solving all the three tasks. Here, we apply an audio-visual pyramid transformer inspired by [8, 15].
@@
Specifically, the tokenized visual $E_{V}$ and audio $E_{A}$ embedding sequences added with its position embeddings $E_{pos}\in\mathbb{R}^{T\times D}$ first pass $L_{1}$ transformer blocks separately for long-form temporal modeling and noise filtering within each modality. Each transformer block [61] regularly consists of a uni-modal multi-head attention (MHA) and a feed-forward network (FFN) with LayerNorm (LN) and residual connections. Afterward, in order to detect modality-aware instances with various lengths, the obtained uni-modal features are fed into an audio-visual pyramid transformer module to integrate informative signals from two modalities at multiple temporal resolutions. The module consists of $L_{2}$ transformer blocks. In each block, 2x downsampling using strided convolutions is first applied. Then, the cross-modal multi-head attention (MHA) is conducted by assigning the current modality as the key and value vectors while another as the query vector, which is followed by FFN and LN layers. We concatenate the audio-enhanced visual feature $F_{l}^{Va}$ and the visual-enhanced audio feature $F_{l}^{Av}$ from the same pyramid level $l$, and finally get an audio-visual feature pyramid.
$Z^{l}=Concat(F_{l}^{Va},F_{l}^{Av}),\;$ (2)
$Z=\{Z^{1},Z^{2},\dots,Z^{L_{2}}\},$ (3)
where $l=[1,L_{2}]$, $F_{l}^{Va},F_{l}^{Av}\in\mathbb{R}^{T_{l}\times D}$, resulting $Z^{l}\in\mathbb{R}^{T_{l}\times 2D}$, and $T_{l}=T/2^{l-1}$ with a downsampling stride $2^{l-1}$.
III-B Task-Specific Experts
@@
Unifying these three tasks under a single framework is inherently challenging as they focus on instances with different characteristics. Inspired by mixture-of-experts (MoE) networks [62, 63, 64], we introduce task-specific experts in our transformer blocks. Unlike previous MoEs that aim to capture modality-specific information (e.g., vision and language), our experts allow the model to learn distinct knowledge for different tasks. Specifically, as shown in the right of Fig. 2, the output feature $F_{l-1},l\in[1,L_{1}+L_{2}]$ from a previous transformer block first passes a shared multi-head attention (MHA) to align information from different tasks. Then, the task-specific information can be captured by switching to different task experts. The task experts compose a Multiway feed-forward network (FFN), where each expert is a standard FFN consisting of two linear layers and an activation.
$F_{l}^{{}^{\prime}}=\mathrm{MHA}(\mathrm{LN}(F_{l-1}))+F_{l-1},\;$ (4)
@@
III-C Prediction Head Design
After the unified audio-visual encoding in Sec. III-A, the output features $\{Z^{l}\}_{l=1}^{L_{2}}$ from the audio-visual feature pyramid will connect to the classification and regression heads to get localization predictions in a single pass.
Unified language-aware classification head. Since the datasets for different tasks have different label vocabularies, it is a straightforward way to use task-specific heads for classification. However, it results in parameter redundancy and limited flexibility due to fixed categories. Here, we propose to unify three task-specific classification heads into one by taking advantage of large pre-trained models. In detail, we treat the instance categories of each dataset as text information and encode them using a pre-trained text encoder. We highlight that the text encoder is from the same pre-trained tri-modal model [21] as the visual and audio encoders used for video tokenization in Sec. III-A. Thus, we can obtain semantically well-aligned embeddings of all three modalities, which provides a strong prior on measuring the similarity between modalities. To add contextual information, prompts are also customized to help specify labels from different tasks. The used prompt templates are “A visual event of {label}.”, “An audio visual event of {label}.”, and “A sound event of {label}.” for TAL, AVEL and SED, respectively. The texts are encoded as $\mathcal{T}=\{\mathcal{T}_{i}\}_{i=1}^{N_{k}}$, where $N_{k}$ is the number of classes of the dataset for the $k$-th task. Then, the encoded texts are linearly projected to a shared embedding space $D^{\prime}$ with the features from the audio-visual feature pyramid. We obtain the normalized text vector $\hat{\mathcal{T}}=\{\hat{\mathcal{T}_{i}}\}_{i=1}^{N_{k}}\in\mathbb{R}^{N_{k}\times D^{\prime}}$ and the normalized cross-modal feature vector $\hat{Z}^{l}\in\mathbb{R}^{T_{l}\times D^{\prime}}$ from each pyramid level to compute the similarities between them.
@@ III-C Prediction Head Design
where $s^{l}\in\mathbb{R}^{T_{l}\times N_{k}}$ indicates the similarities between $N_{k}$ categories and $T_{l}$ temporal segments in the pyramid level $l$ ($l=[1,L_{2}]$), and $\sigma$ is a learnable scaling factor used to adaptively adjust the magnitude of the similarities as in [65]. Then, a sigmoid function is attached to predict the probabilities of $N_{k}$ classes at each moment. Namely, the class with the largest similarity score is the predicted class of a given segment during inference.
Regression head. We simply apply a lightweight regression head for each task. Each regression head as in [15] consists of three layers of 1D convolution attached with layer normalization and ReLU activation. The parameters of the first two convolutional layers are shared among the three heads. The last layer followed by ReLU outputs the distances to the start/end timestamps of an instance at each moment in the pyramid level $l$ if the instance exists as in Eq. 1.
Loss function. Following [8, 15], we apply two losses to train our model in an end-to-end manner. The sigmoid focal loss [66] $\mathcal{L}_{cls}$ and the generalized IoU loss [67] $\mathcal{L}_{reg}$ are used for classification and regression, respectively. The final loss function for each video is denoted as:
$\mathcal{L}=\sum_{t}(\frac{1}{T_{all}}\mathcal{L}_{cls}+\frac{\lambda}{T_{+}}\mathbb{I}_{t}\mathcal{L}_{reg}),$ (7)
where $T_{all}$ is the total snippet number of all pyramid levels. $\mathbb{I}_{t}$ is an indicator function denoting if a timestamp $t$ contains instances. $T_{+}$ is the number of positive snippets that contain instances across all pyramid levels. $\lambda$ is a coefficient that weights the contribution of $\mathcal{L}_{reg}$. We set $\lambda=1$ by default.
III-D Multi-Task Training
We jointly learn TAL, AVEL and SED tasks in a single, unified architecture by multi-task training. The advantage is that the tasks can potentially learn mutually beneficial knowledge from each other. However, the datasets for the tasks vary greatly in size and difficulty, making it very challenging for joint training. For instance, a single epoch of ActivityNet 1.3 [19] for TAL task corresponds to around 15 epochs of DESED [11] for SED task. Besides, all videos in DESED are only 10 seconds while the longest video in ActivityNet 1.3 exceeds 12 minutes. Hence, following the multi-task training method [50], we use a Round-Robin Batch-Level Sampling strategy to sample batches from tasks one-by-one in a cyclical manner, where one iteration consists of each task forwarding a batch and updating parameters in sequence. It effectively trains the model with an equal proportion of individual tasks even though the datasets have various scales (data size and categories), preventing task interference. The Dynamic Stop-and-Go training scheduler [50] is also applied to monitor the validation losses of each task to avoid overfitting.
IV Experiments
IV-A Datasets and Evaluation Metrics
Figure 3: Visualization of dataset statistics. The circle area is proportional to the number of videos in each dataset. There are substantial gaps in dataset scales and video/instance duration between these task-specific datasets.
@@ IV-A Datasets and Evaluation Metrics
There exist significant gaps between these three task-specific datasets in video/instance duration and data scales as shown in Fig. 3, which makes the unification of TAL, AVEL and SED tasks a very challenging problem. Specifically, ActivityNet 1.3 for TAL has many more videos with an average event duration of about 50 seconds, while DESED for SED has fewer videos with events averaging only 1 second. Training on such data naturally leads the TAL task to benefit more from long-duration events, while the SED task relies on short-term signals. We chose these datasets as they are currently the most mainstream datasets with relatively high quality. ActivityNet 1.3 has rich common human activities, which has domain overlaps with UnAV-100 and DESED, facilitating the model to learn mutually beneficial knowledge across tasks during multi-task training. Besides, only UnAV-100 for AVEL is based on untrimmed videos, and DESED is the only dataset whose audio and visual modalities are both available for SED task.
TABLE I: Comparison of our multi-task models to single-task performance. We use both audio and visual modalities for all models. “ST”: single-task, “AT”: all tasks. “All Tasks Average” is computed by averaging the average mAP results of all three tasks. $\cdot\mathrm{M}(\cdot)$”: total $\cdot\mathrm{M}$ parameters needed for $\cdot$ models to conduct all three tasks.
ActivityNet 1.3 (TAL) UnAV-100 (AVEL) DESED (SED)
All Tasks
Average
@@
2 Naive Multi-Task (Base) 55.4 34.9 6.3 34.8 53.0 47.4 41.0 33.5 21.5 49.8 62.7 50.5 26.2 58.6 47.7 62M (1)
3 Ours (TAL & AVEL) 56.4 35.7 5.0 35.7 54.0 48.9 41.9 34.0 21.7 50.8 - - - - - 97M (1)
4 Ours (TAL & SED) 56.5 36.1 4.2 35.5 - - - - - - 61.8 49.3 26.1 58.2 - 97M (1)
5 Ours (AVEL & SED) - - - - 51.8 46.8 40.1 30.0 18.2 48.3 62.6 49.4 27.2 59.4 - 97M (1)
6 OursAT (All-Task) 57.1 36.4 5.5 36.1 54.1 48.6 42.1 34.3 20.5 50.7 64.0 49.5 27.0 59.8 48.9 130M (1)
7 OursAT→ST 56.8 36.0 6.7 36.2 54.8 49.4 43.2 35.3 22.5 51.7 65.1 50.9 26.1 61.1 49.7 186M (3)
IV-B Implementation Details
The sampling rates of sounds and video frames are 16 kHz and 16 fps, respectively. We feed 16 consecutive frames using a sliding window with stride 8 for ActivityNet 1.3, and downsample the features into a fixed length of 256 following [8]. For UnAV-100 and DESED, we use the stride of 4, and pad or crop the feature sequences to 256 and 64, respectively. For each corresponding 1s audio segment, the same stride duration (0.5/0.25s) is used to temporally align with the visual ones. The visual, audio and text encoders of ONE-PEACE [21] are utilized to extract semantically aligned three modality embeddings. The visual encoder is further fine-tuned on Kinetics-400 [60] by freezing the pre-trained parameters and adding several MLP layers, enhancing temporal awareness while maintaining good alignment with other modalities. The extracted feature dimension is 1,536 for all three modalities. In the audio-visual pyramid transformer, the number of attention heads is 4 in both uni-modal and cross-modal blocks. The temporal downsampling operation is realized by using a single depth-wise 1D convolution. The dimension of the embedding space in the framework is $D=D^{{}^{\prime}}=512$, and $L_{1}=2$, $L_{2}=6$. During training, we use Adam for optimization and simply set the same hyperparameters for all datasets of the three tasks. Specifically, the mini-batch size is 16, the initial learning rate is 1e-3 and a cosine learning rate decay is used. Our model is trained only for 5 epochs with a linear warmup of 2 epochs and the weight decay is 1e-4. During inference, our model outputs the on/offsets and classes with confidence scores for all three types of instances occurring in a given video. The output candidates are then processed by Soft-NMS [69] to eliminate highly overlapping ones.
TABLE II: Cross-task transfer performances. Only the head layers of the source-task models are fine-tuned for target tasks during transfer. The results of the single-task models without transfer are in bold.
Source Task Target Task TAL AVEL SED
0.5 Avg. 0.5 Avg. 0.5 Avg.
TAL 56.6 35.3 30.6 29.8 28.1 27.8
AVEL 35.2 21.3 53.3 49.6 43.4 42.9
SED 2.5 1.3 7.4 9.0 61.0 57.7
IV-C Results and Analysis
@@
Single-task v.s. multi-task. To demonstrate the effectiveness of our unified framework, in Table I, we compare our all-task (AT) model (row 6) with two baselines, i.e., single-task (ST) and naive multi-task (Base) in row 1-2. Single-task (ST) models were trained individually on the three tasks, using standard transformer blocks with one FFN and a 1D-conv classifier similar to the regressor. The naive multi-task (Base) model has the same architecture as ST models except for using multi-task training strategies in Section III-D. In other words, the Base model does not use proposed task-specific experts and the language-aware classifier compared with our AT model. We use the same hyperparameters described in Section IV-B as the default setting unless specified otherwise. We observe that naive multi-tasking cannot get decent results on the three tasks (row 2). The performance of the Base model on SED and AVEL saw a slight improvement, due to beneficial knowledge learned from each other. However, the results on TAL declined since there exists a significant gap between ActivityNet 1.3 and other datasets. In contrast, our all-task (AT) model (row 6) that applies the task-specific experts and the unified language-aware classifier can achieve performance boosts on all three tasks compared with ST models, e.g., $0.8\%$, $1.1\%$ and $2.1\%$ at the average mAP on TAL, AVEL and SED, respectively, and $1.4\%$ at the average performance of all tasks. Besides, the total number of parameters reduces by a factor of $1.4\times$ (i.e., 186M to 130M), going from 3 full models to only 1 required for all tasks. It implies both the effectiveness and efficiency of our model.
@@
TABLE III: Comparison with existing state-of-the-art methods. We report the mAP at tIoU=0.5 and the average mAP on three tasks. Best results are in bold and second best underlined. “OP-V/A” denotes the visual/audio encoder of ONE-PEACE [21]. “*” denotes that the results for the AVEL task are from UnAV [15].
@@
Method Visual Encoder Audio Encoder ActivityNet 1.3 (TAL) UnAV-100 (AVEL) DESED (SED)
All Tasks
Average
@@
Cross-task transfer. We perform transfer learning across tasks. We only fine-tune the last layer of the classification and regression heads of the single-task (ST) models during transfer. In Table II, we can see that transfer learning yields very poor results, especially, when transferring the SED single-task model for TAL and AVEL tasks. Besides, cross-task transfer between TAL and AVEL also shows significant performance drops compared to their single-task models, e.g., $19.8\%$ at the average mAP when transferring TAL to AVEL compared to the results of the AVEL ST model. It demonstrates the substantial task/data gap between tasks, thus the unification of the three tasks is a challenging and non-trivial problem.
Pair-wise task relationships. We also explore pair-wise task relationships by jointly training two of three tasks, as shown in rows 3-5 in Table I. We observe that when applying our proposed task-specific experts and unified language-aware classifier, both AVEL and SED (row 3-4) can benefit from the rich common instances in ActivityNet 1.3 when jointly trained with TAL, leading to an obvious improvement compared with the single-task models, i.e., $1.2\%$ and $0.5\%$ at the average mAP, respectively. Besides, SED gains a significant boost when trained with AVEL (row 5), i.e., $1.7\%$ at the average mAP. It could be attributed to category overlap between UnAV-100 and DESED, and training with the large dataset can help prevent overfitting in the small one. Conversely, due to the large gap in instance duration and dataset scales, SED tends to have a negative effect on the AVEL task, resulting in a decrease of $1.3\%$ at the average mAP (row 5). But this effect can be regulated by jointly training all three tasks together using our proposed AT model (row 6).
Multi-task learning as pre-training. Inspired by [50], we fine-tune each single-task model on our trained AT model to demonstrate that the AT model can allow downstream tasks to take advantage of multi-task training. The results are shown in row 7 of Table I, where we initialize ST models using the trained AT model and fine-tune them individually using the same training recipe as in row 1. We can see that the ST models fine-tuned on our AT model outperform the single-task models in row 1 by a large margin, i.e., $0.9\%$, $2.1\%$ and $3.4\%$ on TAL, AVEL and SED, respectively. It indicates that joint training can capture knowledge that is mutually beneficial to all these three tasks, being an effective pre-training step for single-task models.
TABLE IV: Ablation study on the main proposed components. “E-$L_{1}$” denotes applying task-specific experts on the early $L_{1}$ transformer blocks, and “E-$L_{2}$” denotes applying experts on the later $L_{2}$ blocks. “LCH” is short for language-aware classification head. “Prompt” denotes using prompts when tokenizing instance categories for different tasks.